# TDE-3: An improved prior for optical flow computation in spiking neural networks

### v1

@@ 3.3.1 Translating boxes
We start the analysis of TDEs responses to boxes moving to the right (Figure 7) by estimating the direction-selectivity of individual TDEs in a manner similar to DSI (Figure 2). Although strictly speaking, here we cannot calculate the DSI because the motion of the boxes was confined to one direction, we can approximate it by calculating the fraction of total spikes in the network that belonged to the detectors tuned in the L-R direction. Ideally, this number should be one, as only L-R detectors should spike. Table 3 shows that for the network of TDE-2 only 43% of spikes are from the L-R detectors, while for the network of TDE-3 70% of spikes are attributed to L-R detectors. Thus, just as in the case of synthetic data (Figure 2), the TDE-3 has more robust direction-selectivity.
Next, we visualized optical flow coding by the TDEs using the colormap from Figure 12 on Figure 7. Surprisingly, although the TDE-3 has a more robust direction-selectivity on the level of individual detectors, on the network level the TDE-2 and TDE-3 seem to have similar qualitative performance regardless of inference methods. This qualitative observation is also confirmed quantitatively: Table 3 indicates that in all of the testing conditions, the Average Angular Error (AAE) is $\approx\ 20$°) and standard deviation of the angular error ( $\approx\ 30$°) were almost identical.
@@
Though subtraction of velocity estimates made by opposingly tuned detectors mitigates the loss of direction-selectivity of TDE-2 in a textured real-world environment, it is still better to employ TDE-3, because the inhibitory input eliminates residual activity immediately at the level of motion detection. As a result, the network becomes more computationally efficient as it spikes almost two times less (Table 3), and therefore will be more energy-efficient when implemented in neuromorphic hardware [43].
@@ 3.3.2 Rotating disk
Figure 8 employs color-coding scheme from Figure 12 to visualize optical flow coding by various networks of TDEs to compare with ground truth (Figure 8 left bottom). The Figure indicates that in all 4 testing conditions, there is a close correspondence between TDE inference and ground truth. This observation is also supported quantitatively. In all four cases, TDE showed similar (and fairly low, see [39]) average angular error (mean and std of $\approx\ 20$°, Table 4) and relative average endpoint error (mean of $\approx\ 0.4$, std of $\approx\ 0.2$).
Although a 40% error in velocity estimation might look substantial in terms of coding precision, one should keep in mind three things. First of all, we encode over a wide dynamic range of velocities. Secondly, each detector encodes velocities with few spikes/short ISI, so, naturally, such coarse binning would impair precision. However, most importantly, there is almost 90% correlation between TDE velocity estimate and ground truth (Table 4). Indeed, optical flow depends on the ratio between velocity and distance, so it does not provide information about absolute velocity anyway [45]. Meanwhile, high correlation with velocity indicates that regardless of the exact scale, networks of TDE can reliably determine the relative distribution of the velocities in the visual scene.
@@ 4 Discussion
In this work, we made a thorough investigation of the TDE as a computational prior for optical flow estimation in spiking neural networks on three different levels: detector architecture, velocity inference, and real-world performance. Below we summarize the key outcomes of the study and discuss possible future improvements on each of these three levels.
@@
On the level of detector architecture, we showed that classical two-point TDE loses direction-selectivity in a textured environment due to the residual activity in the gain compartment (Figures 1, 2, 7). We proposed a bio-inspired solution for this problem: to augment the TDE-2 with an inhibitory input that resets the gain and removes residual activity upon activation (Figures 1, 2, 7). This procedure preserves the direction-selectivity of individual TDE-3s (Figure 2 and Table 3) and increases the computational efficiency of the network of TDEs by decreasing the total number of spikes almost two-fold. which decreases energy consumption proportionally when implemented on neuromorphic hardware [43].
@@
Although we obtained a principal result about the importance of inhibitory input for the performance of the TDE, there are two issues one needs to address to force TDE-3 to optimal performance. First of all, as was mentioned in Section 5, the timing of events is susceptible to jitter [18, 20]. This creates problem for direction-selectivity as the reset of the gain during motion in the orthogonal direction is due to the simultaneous activation of facilitator and inhibitor compartments (See Section 2 5.2), which is hardly achievable with real-world data due to jitter. A solution to this problem is to low-pass filter inputs to the facilitator and inhibitor compartments. Here it was done crudely by binning events into timesteps (which is still compatible with digital neuromorphic processors such as Loihi [8]), but the TDE could benefit from the incorporation of a proper low-pass filter with trainable parameters.
Along the same lines, inhibition was implemented as instantaneous multiplication by zero. Yet, in an insect elementary motion detector, which serves as inspiration for the TDE-3, inhibition has its dynamics [36]. We think that the TDE-3 would also benefit from inhibition having a prolonged effect on the TDE. For one, it would increase the scope of computations performed by the TDE-3 allowing it to band-pass filter input signals. However, to implement it one needs to carefully think through how corresponding parameters would be trained and in what way (e.g., clipping at zero, suppression with decaying amplitude, etc.) and to which compartment (i.e., not necessarily the gain) this prolonged inhibition should be implemented.
@@
For the velocity inference, we developed a procedure for supervised training of the TDEs to linearly encode input velocities using spike count and ISI-based inference (Figures 9,10). To the best of our knowledge, here we present the first-ever procedure to train spiking neurons to have a particular ISI (Section 5.2). We investigated the performance of both inference methods in conditions with wide (Figure 3) and narrow (Figure 4) dynamic ranges of input velocities and tested their robustness towards variation in stimulus spatial frequency (Figure 5) and background activity noise (Figure 6). We found that although in a noise-free environment ISI-based inference performs slightly better than spike count in terms of the correlation coefficient between true and estimated velocity (Figures 3, 4, 5) when the noise is present spike count turns out to be more reliable, especially when both training and test inference are count based. (Figure 6, Table 2).
@@
Overall, we found that the TDE-based networks are robust in wide range of noise conditions, showing FTA of >95% and correlation with stimulus velocity of $\approx 90\%$ at all noise levels except the two highest (5 Hz/px and 10 Hz/ px, Table 2). However, we suspect that in those conditions it is possible to leverage efficient TDE performance. To this end, we suggest more aggressive filtering at STCF stage, as was done for the real-world data (Figures 7, 8).
@@
In terms of the velocity coding, the fundamental limitation of the TDEs is the latency vs. dynamic range dilemma (see Section 2). Since spike amplitude is usually fixed, high dynamic range requires a lot of spikes (or timepoints between spikes), which necessarily leads to a long latency in velocity inference. The best way to solve this problem is to implement TDE on the Loihi2 neuromorphic chip ([46, 33] that allows "graded" spikes of different amplitude. As a result, one does not need to count spikes or intervals between spikes, because velocity magnitude is immediately available from the first spike. However, apart from Loihi2, no neuromorphic hardware supports spikes of different amplitude. Therefore, one could also solve the latency vs. dynamic range dilemma by using several detectors tuned to different velocity bandwidths such that each detector emits only a few spikes [30].
@@
This was the approach we employed in experiments with real-world data (Figure 8). To the best of our knowledge, these experiments were the first ever to evaluate velocity coding with TDEs in real-world data. Comparison between TDE-2 and TDE-3 performance with the real-world data shows that despite higher energy efficiency of TDE-3 network, both detectors perform almost equally in terms of determination of motion direction (AAE, Table 3, 4) and motion velocity (AEE, Table 3, Table 4) across 360°of change in direction and across the 80-fold change in velocity magnitude (Figure 8). TDEs performed with precision similar to model-based elementary motion detectors [39], but have the advantage of being naturally compatible with spiking neural networks and thus can be implemented in low-power neuromorphic hardware to exploit the spatio-temporal sparsity of event-based streams.
@@
Subtraction of velocity estimates obtained from opposingly tuned detectors improves TDE velocity coding on the network level and seems like a very intuitive post-processing step. Yet, at this level, we still deal with very shallow networks. Future work should aim at the incorporation of TDE-3 into deeper spiking networks with learnable interactions between individual TDE-3 and interplay between spike count and ISI-based inferences. Deeper networks would also allow to mitigate drawbacks (such as latency vs. dynamic range dilemma) and highlight the benefits of the TDE-3.
@@
To conclude, we presented in this paper the TDE-3 as a potent computational prior for motion detection from event-based data in spiking neural networks. It can be trained and allows fast, efficient, precise, and robust coding of optical flow. In a broader sense, our study highlights the importance of understanding the relevant biological features (e.g. relation between structure and function) in the design of neuromorphic solutions for real-time and low power motion detection.
@@ 5.1.1 Synthetic data
The synthetic dataset consists of vertically oriented bars of three intensity values: white, grey, and black. The grayscale values were converted into events using a simplistic custom-made simulator of an event-based camera, which retained its key property: encoding changes in grayscale value with spiking signals that report increases (on events) or decreases (off events) in light intensity. Specifically, we modeled voltage in the event-based detector as being proportional to the logarithm of light intensity and reported contrast each time the temporal derivative of this signal $(V_{t}-V_{t-1})$ was higher/lower than the threshold value of ±$\pm 0.15$.
To compare the robustness in direction-selectivity of TDE-3 and TDE-2, we simulated their responses to moving textured stimuli consisting of bars oriented orthogonally towards the direction of motion. The bars had three intensity levels: white, grey, and black. The size of the texture along the motion axis was 80 pixels and 3 pixels along the orthogonal direction (Figure 2A). We employed 5 velocities: 0.1 px/timestep, 0.2 px/timestep, 0.33 px/timestep, 0.5 px/timestep and 1 px/timestep. The motion direction (left-right, right-left, top-bottom, bottom-top) and velocity were randomly chosen for each stimulus example (2000 per epoch). To vary the "amount" of texture in stimuli we randomly varied the fraction of the grey bars from 0% to 80%.
@@
To study the effect of noise on training and inference with spike count and ISI-based methods, we simulated one of the most prominent types of noise in event-based vision - background activity noise, as a Poisson process following the approach by [16]. The noise was injected after the conversion of grayscale values to events as some stochastic events. We used six mean rates of noise: 0.2 Hz/px, 0.5 Hz/px, 1 Hz/px, 2 Hz/px, 5 Hz/px, and 10 Hz/px. This effectively covered the dynamic range in event-based vision from high to low SNR conditions [16]. For this set of experiments, we employed the same set of 5 velocities as above and varied the distance between edges within a range of 3 to 10 pixels as above. The timestep duration was set to 10ms.

### v2

@@ 3.3.1 Translating boxes
We start the analysis of TDE’s responses to boxes moving to the right (Figure 10) by estimating the direction-selectivity of individual TDEs in a manner similar to DSI (Figure 2). Although strictly speaking, here we cannot calculate the DSI because the motion of the boxes was confined to one direction, we can approximate it by calculating the fraction of total spikes in the network that belonged to the detectors tuned in the L-R direction. Ideally, this number should be one, as only L-R detectors should spike. Table 2 shows that for the network of TDE-2 only 43% of spikes are from the L-R detectors, while for the network of TDE-3 70% of spikes are attributed to L-R detectors. Thus, just as in the case of synthetic data (Figure 2), the TDE-3 has more robust direction-selectivity.
Next, we visualized optical flow coding by the TDEs using the colormap from Figure 9 on Figure 10. Surprisingly, although the TDE-3 has a more robust direction-selectivity on the level of individual detectors, on the network level the TDE-2 and TDE-3 seem to have similar qualitative performance regardless of inference methods. This qualitative observation is also confirmed quantitatively: Table 2 indicates that in all of the testing conditions, the Average Angular Error (AAE) ( $\approx\ 20$°) and standard deviation of the angular error ( $\approx\ 30$°) were almost identical.
@@
Though subtraction of velocity estimates made by opposingly tuned detectors mitigates the loss of direction-selectivity of TDE-2 in a textured real-world environment, it is still better to employ TDE-3, because the inhibitory input eliminates residual activity immediately at the level of motion detection. As a result, the network becomes more computationally efficient as it spikes almost two times less (Table 2), and therefore will be more energy-efficient when implemented in neuromorphic hardware [52]. To make a statistical argument for the higher energy efficiency of TDE-3 we also compared detector spike count with an additional 10 real-world scenes (rotating disk + 9 scenes from [56]) and found that in all of the visual scenes, TDE-2 spiked more than TDE-2 (on average $2.6\pm 0.75$. times,p=1e-4; see Supplementary Information Figure S2 for the details).
@@ 3.3.2 Rotating disk
Figure 11 employs a color-coding scheme from Figure 9 to visualize optical flow coding by various networks of TDEs to compare with ground truth (Figure 11 left bottom). The Figure indicates that in all 4 testing conditions, there is a close correspondence between TDE inference and ground truth. This observation is also supported quantitatively. In all four cases, TDE showed similar (and fairly low, see [48]) AAE (mean and std of $\approx\ 20$°, Table 3) and relative average endpoint error (rAEE, mean of $\approx\ 0.4$, std of $\approx\ 0.2$).
Although a 40% error in velocity estimation might look substantial in terms of coding precision, one should keep in mind three things. First of all, we encode over a wide dynamic range of velocities. Secondly, each detector encodes velocities with few spikes/short ISI, so, naturally, such coarse binning would impair precision. However, most importantly, there is almost 90% correlation between TDE velocity estimate and ground truth (Table 3). Indeed, optical flow depends on the ratio between velocity and distance, so it does not provide information about absolute velocity anyway [57]. Meanwhile, a high correlation with velocity indicates that regardless of the exact scale, networks of TDE can reliably determine the relative distribution of the velocities in the visual scene.
@@ 4 Discussion
In this work, we made a thorough investigation of the TDE as a computational prior for optical flow estimation in SNNs on three different levels: detector architecture, velocity inference, and real-world performance. Below we summarize the key outcomes of the study and discuss possible future improvements on each of these three levels.
4.1 Detector architecture
@@
On the level of detector architecture, we showed that classical TDE-2 loses direction-selectivity in a textured environment due to the residual activity in the gain compartment (Figures 1, 2, 10). We proposed a bio-inspired solution for this problem: to augment the TDE-2 with an inhibitory input that resets the gain and removes residual activity upon activation (Figures 1, 2, 10). This procedure preserves the direction-selectivity of individual TDE-3s (Figure 2 and Table 2) and increases the computational efficiency of the network of TDEs by decreasing the total number of spikes 2-4 times (Supplementary Figure S2), which decreases energy consumption proportionally when implemented on neuromorphic hardware [52].
Although we obtained a principal result about the importance of inhibitory input for the performance of the TDE, there are two issues one needs to address to force TDE-3 to optimal performance: jitter in event timing [20, 22] and the way inhibition is implemented.
@@
The jitter creates a problem for direction-selectivity because the reset of the gain during motion in the orthogonal direction is due to the simultaneous activation of facilitator and inhibitor compartments (See Sections 2, 5.2), which is hardly achievable with real-world data due to the jitter. A solution to this problem is to low-pass filter inputs to the facilitator and inhibitor compartments. Here it was done crudely by binning events into timesteps (which is still compatible with digital neuromorphic processors such as Loihi [8]), but the TDE could benefit from the incorporation of an adaptive low-pass filter with trainable parameters. It is crucial for the filter to be adaptive because there is a trade-off between velocity resolution that requires short time constants and combating the noise that requires long time constants. Adaptability is crucial since the optimal time constant depends on the particular characteristics of the visual scene such as velocity distribution, illumination, contrast, spatial frequency, etc.
In the present study, inhibition was implemented as an instantaneous reset of the gain, modeled as a multiplication by zero. In contrast, in the insect elementary motion detector that inspired TDE-3, inhibition has its own dynamics and can influence neuronal resistance, current, membrane potential, and time constant [54]. These richer dynamics have mainly been studied in detailed biophysical models of insect vision [54].
Notably, nearly two decades before 3-point elementary motion detectors were discovered in insects, a similar triplet architecture was proposed for neuromorphic hardware by Kramer [37]. His FTI (facilitator-trigger-inhibitor) detector, as well as related approaches [58, 59, 60], used inhibition to suppress spiking by acting directly at the output. While TDE-3 shares the same functional goal — suppressing responses to motion in the non-preferred direction — it implements inhibition by resetting the gain, thereby intervening earlier in the computation. Other applications of triplet correlations for optical flow estimation from event-based data [61] used simple averaging of velocity estimates between triplet pairs and did not rely on spiking neural networks or trainable systems. Grayscale-based detectors [62, 63] modeled inhibition as a linear subtraction between the activations of the inhibitor and facilitator compartments. These diverse implementations highlight the flexibility of triplet-based motion detection architectures.
In the present study, inhibition was implemented as an instantaneous multiplication of the gain by zero. However, in the insect elementary motion detector, which serves as inspiration for TDE-3, inhibition has its own dynamics and influences neuronal resistance, current, membrane potential, and time constant [54]. Nevertheless, such complex interactions have primarily been studied in the context of detailed biophysical models of insect motion detection [54]. Notably, nearly two decades before 3-point elementary motion detectors were discovered in insects, similar architectures had already been proposed for neuromorphic hardware by Kramer [37] from a purely engineering perspective. In his FTI (facilitator-trigger-inhibitor) detector, as well as in related approaches [58, 59, 60], the inhibitor acted directly on the neuronal output, preventing firing upon activation. Other applications of triplet correlations for optical flow estimation from event-based data [61] employed a simple averaging of velocity estimates between each pair within the event triplet and did not use SNNs or any type of neural network. Studies using grayscale images [62, 63] implemented detectors where inhibition was modeled as a simple subtraction of the inhibitor’s activation from that of the facilitator.
Future research could benefit from exploring more biologically accurate inhibition dynamics and systematically investigating different interaction modes between TDE-3 compartments.
4.2 Velocity inference
For the velocity inference, we developed a procedure for supervised training of the TDEs to linearly encode input velocities using spike count and ISI-based inference (Figures 12,13). For the ISI-based training, we developed a novel approach. While most other methods [45, 46, 47] model the relationship between input magnitude with differentiable function than use BPTT, we measure ISI from the amplitude of low-pass filtered spike train to use BPTT with surrogate gradient (see Section 5.2.2 for the details). The theoretical advantage of our method is that it does not require estimation of the relationship between spike timing and input magnitude, allowing work with more conventional activation functions.
@@ 4.2 Velocity inference
We investigated the performance of both inference methods in conditions with wide (Figure 3) and narrow (Figure 4) dynamic ranges of input velocities and tested their robustness towards variation in stimulus spatial frequency (Figure 5) and background activity noise (Figure 6). We found that although in a noise-free environment, ISI-based inference performs slightly better than spike count in terms of the correlation coefficient between true and estimated velocity (Figures 3, 4, 5) when the noise is present spike count turns out to be more reliable, especially when both training and test inference are count based (Figures 6, 7).
@@
Overall, we found that the TDE-based networks are robust in a wide range of noise conditions, showing FTA of >95% and correlation with stimulus velocity of $\approx 90\%$ at all noise levels except the two highest (5 Hz/px and 10 Hz/ px, Figure 7). However, we suspect that in those conditions it is possible to leverage efficient TDE performance. To this end, we suggest more aggressive filtering at the STCF stage, as was done for the real-world data (Figures 10, 11).
Examiming training progression we observe that while the loss curves for training with spike count inference tend to be relatively smooth, the loss curves for training with ISI exhibit noticeable jumps, with changes in the loss value appearing more pulsatile than gradual. This behavior can be attributed to the initialization, which results in very high spiking rates (see Section 5.2 for the details). Under these conditions, it is easier to regulate spike count than the ISI (inter-spike interval) between the first two spikes. For spike count, improving the loss function requires only a slight adjustment in the spike count for one of the stimulus velocities. In contrast, ISI regulation is more challenging because the ISI tends to decrease over time (being shortest immediately after trigger activation). Consequently, training with ISI requires more epochs to produce meaningful differences in velocity estimation that also manifest themselves in more abrupt fashion.
To decode stimulus velocity from TDE activity in the presence of multiple edges one needs to perform motion segmentation to relate counted spikes/measured ISIs to a particular moving edge. How to know when to count spikes and to measure ISI? Here we proposed to count spikes starting from each activation of TDE’s trigger as it indicates passage of a new edge. This worked for a shallow network where we could read out trigger activations, but suppose there would be more layers after the TDE. How could downstream circuitry perform motion segmentation? One obvious solution would be to establish so-called skip connections between input to TDE’s trigger and downstream circuitry. Another solution could exploit the fact that immediately after activation of the trigger the TDE’s signal is maximum i.e. if one would count spikes within a certain sliding time window, the local (in time) maximal count would be close to the moment of edge appearance at TDE’s trigger. Similarly, the shortest distance between two spikes would be for the first two spikes after trigger activation. Hence, downstream circuitry potentially can perform motion segmentation by performing temporal max pooling.
@@
For the velocity coding, the fundamental limitation of the TDEs is the latency vs. dynamic range dilemma (see Section 2). Since spike amplitude is usually fixed, high dynamic range requires a lot of spikes (or timepoints between spikes), which necessarily leads to a long latency in velocity inference. The best way to solve this problem is to implement TDE on the neuromorphic chip Loihi2 ([64, 35] that allows "graded" spikes of different amplitude. As a result, one does not need to count spikes or intervals between spikes, because velocity magnitude is immediately available from the first spike. However, apart from Loihi2, no neuromorphic hardware supports spikes of different amplitude. Therefore, one could also solve the latency vs. dynamic range dilemma by employing the so-called "eccentric downsampling" using several detectors tuned to different velocity bandwidths such that each detector emits only a few spikes (Figure 9 [32]).
4.3 Performance with real-world data
@@
To the best of our knowledge, our experiments on Figures 10,11 are the first to evaluate velocity coding with TDEs in real-world data. [^2] A preprint by Greatorex et al. [40] presenting an ego-motion estimation with a network of TDE-2 appeared on arXiv shortly after the initial release of our manuscript. Comparison between TDE-2 and TDE-3 performance with the real-world data shows that despite higher energy efficiency of TDE-3 network (see Table 2, Supplementary Figure S2), both detectors perform almost equally in terms of determination of motion direction (AAE, Table 2, 3) and motion velocity (AEE, Table 2) across 360°of change in direction and the 80-fold change in velocity magnitude (Figure 11). TDEs performed with precision similar to model-based elementary motion detectors [48], but have the advantage of being naturally compatible with SNNs and thus can be implemented in low-power neuromorphic hardware to exploit the spatio-temporal sparsity of event-based streams.
@@
Subtraction of velocity estimates obtained from opposingly tuned detectors improves TDE velocity coding on the network level and seems like a very intuitive post-processing step. Yet, at this level, we still deal with very shallow networks. Future work should aim at the incorporation of TDE-3 into deeper spiking networks with learnable interactions between individual TDE-3 and the interplay between spike count and ISI-based inferences. Deeper networks would also allow one to mitigate drawbacks (such as latency vs. dynamic range dilemma) and highlight the benefits of the TDE-3.
Our experiments with real-world data suggest the training of the TDEs is generalizable. Indeed, TDE-3 was trained in simulation and yet worked fine in the real world. Moreover, we trained a single detector, expressing velocities in abstract quantities of velocities, expressing them as a fraction of the distance between facilitator and trigger covered in one algorithmic timestep (see Section 5.2.3) and subsequently scaled the dynamic range of encoded velocities by simply changing the distance between facilitator and trigger. Thus, due to TDEs simplicity, one can easily adapt it to a variety of conditions with an appropriate choice of spacing between TDE outputs and the duration of the timestep.
We hypothesize that TDE-3 can enhance the generalizability of SNNs by extracting simple, abstract features—such as moving edges—that are ubiquitous in visual scenes. In support of this, research on drone acrobatics has shown that autopilots trained in abstract representations of visual scenes bridge the simulation-to-reality gap more effectively than those trained on raw images [65]. We believe that for low-level visual processing, TDE-3 offers greater robustness compared to Convolutional Neural Networks (CNNs). Unlike CNNs, which rely on architectures shaped by training data and are therefore prone to overfitting, TDE-3 employs a fixed, analytic architecture, reducing the risk of overfitting and enhancing stability.
@@
To conclude, we presented in this paper the TDE-3 as a potent computational prior for motion detection from event-based data in SNNs. It can be trained and allows fast, efficient, precise, and robust coding of optical flow. In a broader sense, our study highlights the importance of understanding the relevant biological features (e.g. relation between structure and function) in the design of neuromorphic solutions for real-time and low-power motion detection.
@@ 5.1.1 Synthetic data
The synthetic dataset consists of vertically oriented bars of three light intensity values: white, grey, and black. The greyscale values were converted into events using a simplistic custom-made simulator of an event-based camera, which retained its key property: encoding changes in greyscale value with spiking signals that report increases (on events) or decreases (off events) in light intensity. Specifically, we modeled voltage in the event-based detector as being proportional to the logarithm of light intensity and reported contrast each time the temporal derivative of this signal $(V_{t}-V_{t-1})$ was higher/lower than the threshold value of ±$\pm 0.15$.
To compare the robustness in direction-selectivity of TDE-3 and TDE-2, we simulated their responses to moving textured stimuli consisting of bars oriented orthogonally towards the direction of motion. The bars had three intensity levels: white, grey, and black. The size of the texture along the motion axis was 80 pixels and 3 pixels along the orthogonal direction (Figure 2A). We employed 5 velocities: 0.1 px/timestep, 0.2 px/timestep, 0.33 px/timestep, 0.5 px/timestep and 1 px/timestep. The motion direction (left-right, right-left, top-bottom, bottom-top) and velocity were randomly chosen for each stimulus example (2000 per testing round). To vary the "amount" of texture in stimuli we randomly varied the fraction of the grey bars from 0% to 80%.
@@
To study the effect of noise on training and inference with spike count and ISI-based methods, we simulated one of the most prominent types of noise in event-based vision - background activity noise, as a Poisson process following the approach by [18]. The noise was injected after the conversion of greyscale values to events as some stochastic events. We used six mean rates of noise: 0.2 Hz/px, 0.5 Hz/px, 1 Hz/px, 2 Hz/px, 5 Hz/px, and 10 Hz/px. This effectively covered the dynamic range in event-based vision from high to low SNR conditions [18]. For this set of experiments, we employed the same set of 5 velocities as above and varied the distance between edges within a range of 3 to 10 pixels as above. The timestep duration was set to 10ms.